From a pack of $52$ cards,two cards are drawn one by one without replacement. The probability that the first drawn card is a king and the second is a queen,is

  • A
    $\frac{2}{13}$
  • B
    $\frac{8}{663}$
  • C
    $\frac{4}{663}$
  • D
    $\frac{103}{663}$

Explore More

Similar Questions

$A$ pandemic has been spreading all over the world. The probabilities are $0.7$ that there will be a lockdown,$0.8$ that the pandemic is controlled in one month if there is a lockdown,and $0.3$ that it is controlled in one month if there is no lockdown. The probability that the pandemic will be controlled in one month is

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

The probability that India wins the toss is $3/4$. If it wins the toss,the probability of winning the match is $4/5$,otherwise it is $1/2$. Find the probability that India wins the match.

Difficult
View Solution

Two cards are drawn from a pack of well-shuffled $52$ playing cards one by one without replacement. The probability that both cards are queens is:

Two cards are drawn at random and without replacement from a pack of $52$ playing cards. Find the probability that both the cards are black.

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo